A beam of neutrons with a de Broglie wavelength of 0.231 nm diffracts from a crystal of table salt, which has an interionic spacing of 0.282 nm

What is the angle of the first interference maximum? __________ degrees
Answer #1

For Bragg scattering, ? = arcsin(m*lambda/(2d)) = 24 deg
Answer #2

For Bragg scattering,

? = arcsin(m*lambda/(2d))

= 0.4454 rad

=25.52 deg
Answer #3

From the   bragg's diffraction formula

2 d sin theat = n L

2(0.282 nm) sin theta = 1(0.231 nm )

theat = sin ^-1 (1(0.231 nm )/2(0.282 nm)) = 24.17 degree
Answer #4

n = 1

we have d*sin(A) = n*lamda

0.282*10^-9*sin(A)= 1*0.231*10^-9

sin(A) = 0.231/0.282

A = 54.999 degrees
Answer #5

i think that 0.125
Answer #6

i think that 24.79
Answer #7

A beam of neutrons with a de Broglie wavelength of 0.231 nm diffracts from a crystal of table salt, which has an interionic spacing of 0.282 nm

What is the angle of the first interference maximum?

60 degrees